# Evaluating Vision–Language and Large Language Models for Automated Student Assessment in Indonesian Classrooms

### v1

@@ Abstract
Although vision-language and large language models (VLM and LLM) offer promising opportunities for AI-driven educational assessment, their effectiveness in real-world classroom settings, particularly in underrepresented educational contexts, remains underexplored. In this study, we evaluated the performance of a state-of-the-art VLM and several LLMs on 646 handwritten exam responses from grade 4 students in six Indonesian schools, covering two subjects: Mathematics and English. These sheets contain more than 14K student answers that span multiple choice, short answer, and essay questions. Assessment tasks include grading these responses and generating personalized feedback. Our findings show that the VLM often struggles to accurately recognize student handwriting, leading to error propagation in downstream LLM grading. Nevertheless, LLM-generated feedback retains some utility, even when derived from imperfect input, although limitations in personalization and contextual relevance persist.
@@ 1 Introduction
However, most VLM and LLM-based educational tools have been developed with English-speaking contexts in mind Lee and Zhai (2025); Yancey et al. (2023), limiting their relevance and usability in non-English-speaking regions, particularly in rural areas in Indonesia. Ensuring socio-cultural relevance is essential: effective deployment requires adaptation to local curricula, languages, and cultural norms, rather than relying on a one-size-fits-all approach. Moreover, the shortage of qualified teachers in rural areas highlights the importance of prioritizing AI integration in underserved regions, rather than concentrating development efforts solely in high-resource, Global North contexts Jin et al. (2025); Kristiawan et al. (2024).
In this study, we address the contextual challenges of applying AI-powered assessment tools in non-English speaking and under-resourced settings by collecting real-world student assessment data from primary schools in the form of handwritten responses. This design is motivated by two practical considerations. First, many schools, especially in rural areas, lack consistent access to digital devices, highlighting the need for AI systems that function effectively in low-tech environments. Second, using handwritten responses helps reduce the risk of academic dishonesty, such as students who rely on AI tools to generate answers. Assessments were conducted in Indonesian for the mathematics subject, while the responses to the English subject were written in English, reflecting the language of instruction for each subject.
Our contributions are as follows: (1) We release a dataset of 646 handwritten student answer sheets (with over 14K answers) collected from six primary schools in Indonesia—three from rural areas and three from urban areas. The assessments cover Grade 4 Mathematics and English, with questions and scoring guidelines developed by experienced teachers. All student responses were manually transcribed and graded by professional teachers. [^1] To ensure ethical use and protect student privacy, all personally identifiable information (e.g., student names, grade levels, and school names) has been removed. (2) We introduce a multimodal pipeline that integrates vision–language models (VLMs) and large language models (LLMs), as illustrated in Figure 1. We compare several state-of-the-art models for grading student answers and find that GPT-4o with vision input achieves the highest accuracy and feedback quality. (3) We conduct a manual evaluation of LLM-generated feedback in Indonesian and find that, even when based on imperfect input (e.g., OCR errors), the feedback tends to be clear and factually correct. However, personalization and helpfulness remain notable areas of concern.
@@ 3 Dataset Construction
We developed assessment instruments for grade 4 primary school students in two subjects: Mathematics and English. The items were designed from scratch based on a thorough analysis of the national curriculum and corresponding learning objectives. Each subject assessment consisted of 10 multiple-choice questions (MCQs), 10 short-answer questions, and 2 essay questions. All items were created by experienced senior subject teachers—an English teacher and a Math teacher—each with over 10 years of classroom experience and a Master’s degree in Education. In addition to writing the assessment items, these teachers developed detailed scoring rubrics for the short-answer and essay questions, as well as answer keys for the MCQs. Standardized answer sheets were also prepared to collect student responses.
@@
Data collection was carried out in six primary schools, evenly divided between rural (Sumatra and Nusa Tenggara Islands) and urban (Java Island) settings. Each classroom included approximately 20 to 30 students. For both subjects, students followed a structured sequence consisting of a pre-test, lesson, and post-test. Students had up to 30 minutes to complete their answers on a standardized answer sheet.
In total, we collected 646 handwritten answer sheets from these assessments, comprising both pre-tests and post-tests. Of these, 414 were collected from urban schools and 232 from rural schools. The disparity in sample size between urban and rural areas is primarily due to larger class sizes typically found in urban schools compared to their rural counterparts.
4 Experiment
Overall Pipeline
Model
For OCR, we use GPT-4o OpenAI et al. (2024), alongside a gold-standard transcription manually parsed by teachers. For automatic scoring, we compare the performance of GPT-4o, Llama-3.1-Instruct (70B) Touvron et al. (2023b), Qwen2.5-Instruct (72B) Team (2024), and Deepseek-Chat (671B) Liu et al. (2024a). For generating personalized feedback, we rely on the scoring results produced by GPT-4o and generate two versions of feedback using GPT-4o and Deepseek-Chat. All prompts and decoding hyperparameters used are provided in the Appendix.
Evaluation
Each answer sheet image was manually transcribed and scored by professional teachers. We compared the LLM-generated scores against these gold-standard scores across three question types: multiple-choice, short-answer, and essay, using mean absolute error (MAE) as the evaluation metric. For personalized feedback, we conducted a manual evaluation covering four aspects—Correctness, Personalization, Clarity, and Educational Value/Helpfulness—rated on a 1–5 scale, where 1 indicates the lowest quality. [^2] This evaluation was carried out by an experienced educator with a Master’s degree in teaching. The evaluation guidelines and definitions for each aspect are provided in the Appendix.
@@ 4 Experiment
Table 1: Mean absolute error (MAE) for scoring English and Math, calculated separately for multiple-choice (M), short-answer (S), essay (E), and the total score. Lower values indicate better performance; bolded numbers represent the best results. Scores for each component range from 0 to 100.
@@ 5 Result and Analysis
Table 1 presents the performance of the LLMs selected in three types of questions: multiple choice, short answer, and essay. When using GPT-4o to extract student responses via OCR, we observe that most model-generated scores are generally competitive. Among them, GPT-4o produces scores that align most closely with human grading for essay questions, achieving the lowest MAE in both English (5.6) and Math (1.5). In contrast, LLaMA-3.1–70B and Qwen-2.5–72B are less reliable, with scores deviating more significantly from human judgments. Short-answer questions remain the most challenging to evaluate: even the best performing model in this category, LLaMA-3.1-7B for Math, still shows a relatively high MAE of 10.6, indicating a notable gap from human-level accuracy.
However, the results differ when human effort is involved in the OCR task. Most scores become better overall, with Deepseek-chat and GPT-4o emerging as the top-performing models. Deepseek-chat shows strong performance in English (MAE of 4.4 for short answers and 1.5 for essays), while GPT-4o performs best in Math, with only a 2.9 difference in short answers and 5.7 in essays. It is worth noting that MCQ scores remain at 0, as basic string matching is sufficient due to the exact nature of the answers. The impact of OCR performance on LLM scoring is further discussed in Section 5.
@@
Table 3: Analysis of mean absolute errors (MAE) for scoring English and Math across urban and rural settings, calculated separately for multiple-choice (M), short-answer (S), essay (E), and total scores. The OCR results used in this analysis were obtained through human transcription. Lower values indicate better performance; bolded values represent the best results. Each component is scored on a 0–100 scale.
@@
Given the differences in MAE between the GPT-4o OCR outputs and human transcription shown in Table 1, we further analyze the OCR performance of GPT-4o and evaluate the extent to which recognition errors propagate to the subsequent scoring. For this analysis, we use exact string matching to assess accuracy on multiple choice and short answer questions, and compute ROUGE-L Lin (2004) scores to compare GPT-4o and human transcriptions for essay questions.
@@ 6 Conclusion
In this work, we present a real-world implementation of vision–language model (VLM) and large language models (LLMs) for student assessment in underrepresented regions—specifically, rural and urban areas of Indonesia—focusing on primary school subjects in Math and English. Our results show that GPT-4o and Deepseek (671B) perform competitively in matching teacher-assigned scores across multiple-choice, short-answer, and essay formats. For personalized feedback generation, manual evaluation indicates that Deepseek outperforms GPT-4o in terms of quality and relevance. We hope that this work encourages greater research attention towards educational applications of AI in low-resource and underserved contexts.
@@ Acknowledgments
This research was supported by the 2023 SEAMEO-Australia Education Links Award, jointly organized by the Southeast Asian Ministers of Education Organization (SEAMEO) Secretariat and the Australian Government Department of Education.
@@
Appendix A Hyperarameter Setup
@@
Appendix B Prompts List
@@
Appendix C Human Evaluation Guideline on Personalised Feedback

### v2

# From Handwriting to Feedback: Evaluating VLMs and LLMs for AI-Powered Assessment in Indonesian Classrooms

@@ Abstract
Despite rapid progress in vision–language and large language models (VLMs and LLMs), their effectiveness for AI-driven educational assessment in real-world, underrepresented classrooms remains largely unexplored. We evaluate state-of-the-art VLMs and LLMs on over 14K handwritten answers from grade-4 classrooms in Indonesia, covering Mathematics and English aligned with the local national curriculum. Unlike prior work on clean digital text, our dataset features naturally curly, diverse handwriting from real classrooms, posing realistic visual and linguistic challenges. Assessment tasks include grading and generating personalized Indonesian feedback guided by rubric-based evaluation. Results show that the VLM struggles with handwriting recognition, causing error propagation in LLM grading, yet LLM feedback remains pedagogically useful despite imperfect visual inputs, revealing limits in personalization and contextual relevance.
@@ 1 Introduction
However, most VLM- and LLM-based educational tools have been developed in Global North contexts Lee and Zhai (2025); Yancey et al. (2023), limiting their applicability in classrooms where curricula, languages, and instructional practices differ substantially. Even within these well-resourced settings, only a few studies (e.g., Daniela et al. 2025) have explored end-to-end AI assessment on handwritten student work—and typically on very small samples, such as 24 responses—while the majority of existing benchmarks continue to rely on clean, digital text Chiang et al. (2024).
In Indonesia, paper-based assignments remain the preferred practice in most schools. Many schools, especially in rural areas, lack consistent access to digital devices, highlighting the need for AI systems that function effectively in low-tech environments. Second, using handwritten responses helps reduce the risk of academic dishonesty, such as students who rely on AI tools to generate answers.
Our contributions are as follows:
  • We release the first dataset of handwritten student answer sheets for AI-based educational assessment, comprising over 14K responses collected from six primary schools in Indonesia—three rural and three urban. The dataset covers Grade 4 Mathematics and English, with questions and scoring rubrics designed by experienced teachers. All answers were manually transcribed and graded by professional educators. [^1] To ensure ethical use and protect student privacy, all personally identifiable information (e.g., student names, grade levels, and school names) has been removed. In Indonesia, English is taught as a second language, while Mathematics often involves handwritten explanations in Indonesian (see essay examples in Figure 5).
  • We introduce a multimodal pipeline that integrates vision–language models (VLMs) and large language models (LLMs), as illustrated in Figure 1. We compare several state-of-the-art models for grading student answers and find that GPT-4o with vision input achieves the highest accuracy and feedback quality.
  • We conduct a manual evaluation of LLM-generated feedback in Indonesian and find that, even when based on imperfect input (e.g., OCR errors), the feedback tends to be clear and factually correct. However, personalization and helpfulness remain notable areas of concern.
@@ 3 Dataset Construction
We developed assessment instruments for grade 4 primary school students in two subjects: Mathematics and English. The items were designed from scratch based on a thorough analysis of the national curriculum and corresponding learning objectives. Each subject assessment consisted of 10 multiple-choice questions (MCQs), 10 short-answer questions, and 2 essay questions. All items were created by experienced senior subject teachers—an English teacher and a Math teacher—each with over 10 years of classroom experience and a Master’s degree in Education. In addition to writing the assessment items, these teachers developed detailed scoring rubrics for the short-answer and essay questions, as well as answer keys for the MCQs. Standardized answer sheets were also prepared to collect student responses. The assessment and rubric examples for both English and Math subjects are presented in Appendix A.
@@
Data collection was carried out in six primary schools, evenly divided between rural (Sumatra and Nusa Tenggara Islands) and urban (Java Island) settings. Each classroom included approximately 20 to 30 students. Students had up to 30 minutes to complete their answers on a standardized answer sheet.
In total, we collected 646 handwritten answer sheets from these assessments, with a total of more than 14K student answers. Of these, 414 were collected from urban schools and 232 from rural schools (See Figure 5 and Figure 6 in Appendix for example of student answer sheets). The disparity in sample size between urban and rural areas is primarily due to larger class sizes typically found in urban schools compared to their rural counterparts.
@@
Table 1: Mean absolute error (MAE) for scoring English and Math, calculated separately for multiple-choice (M), short-answer (S), essay (E), and the total score. Lower values indicate better performance; bolded numbers represent the best results. Scores range from 0 to 100.
4 Experiment
Overall Pipeline
Model
For OCR, we employ GPT-4o OpenAI et al. (2024), alongside a gold-standard transcription manually prepared by teachers for comaprison. We also evaluated Qwen-VL-72B (see Table 5 in Appendix E) but found its performance suboptimal; hence, GPT-4o was used for all subsequent experiments. For automatic scoring, we compare the performance of GPT-4o, Llama-3.1-Instruct (70B) Touvron et al. (2023b), Qwen2.5-Instruct (72B) Team (2024), and DeepSeek-Chat (671B) Liu et al. (2024a). For personalized feedback generation, we rely on the grading results from GPT-4o, producing two feedback variants using GPT-4o and DeepSeek-Chat, respectively. All prompts and decoding hyperparameters are detailed in Appendix B and Appendix C.
Evaluation
Each answer sheet image was manually transcribed and scored by professional teachers. We compared the LLM-generated scores against these gold-standard scores across three question types: multiple-choice, short-answer, and essay, using mean absolute error (MAE) as the evaluation metric. For personalized feedback, we conducted a manual evaluation covering four aspects—Correctness, Personalization, Clarity, and Educational Value/Helpfulness—rated on a 1–5 scale, where 1 indicates the lowest quality. [^2] This evaluation was carried out by experienced educators with a Master’s degree in teaching. The evaluation guidelines and definitions for each aspect are provided in the Appendix.
@@ 5 Result and Analysis
Table 1 presents the performance of the LLMs selected in three types of questions: multiple choice, short answer, and essay. When using GPT-4o to extract student responses via OCR, we observe that most model-generated scores are generally competitive. Among them, GPT-4o produces scores that align most closely with human grading for essay questions, achieving the lowest MAE in both English (5.6) and Math (1.5). In contrast, LLaMA-3.1–70B and Qwen-2.5–72B are less reliable, with scores deviating more significantly from human judgments. Short-answer questions remain the most challenging to evaluate: even the best-performing model in this category, LLaMA-3.1-7B for Math, still shows a relatively high MAE of 10.6, indicating a notable gap from human-level accuracy.
However, the results differ when human effort is involved in the OCR task. Most scores become better overall, with Deepseek-chat and GPT-4o emerging as the top-performing models. Deepseek-chat shows strong performance in English (MAE of 4.4 for short answers and 1.5 for essays), while GPT-4o performs best in Math, with only a 2.9 difference in short answers and 5.7 in essays. It is worth noting that MCQ scores remain at 0, as basic string matching is sufficient due to the exact nature of the answers. The impact of OCR performance on LLM scoring is further discussed in Section 5 of OCR performance analysis.
@@
Table 3: MAE for scoring English and Math across urban and rural settings, calculated separately for multiple-choice (M), short-answer (S), essay (E), and total scores. The OCR results used in this analysis were obtained through human transcription. Lower values indicate better performance; bolded values represent the best results. Each component is scored on a 0–100 scale.
@@
Given the differences in MAE between the GPT-4o OCR outputs and human transcription shown in Table 1, we further analyze the OCR performance of GPT-4o and evaluate the extent to which recognition errors propagate to the subsequent scoring. For this analysis, we use exact string matching to assess accuracy on multiple-choice and short answer questions, and compute ROUGE-L Lin (2004) scores to compare GPT-4o and human transcriptions for essay questions.
@@ 6 Conclusion
We release the first dataset of handwritten student answer sheets for AI-based assessment in Indonesia, covering over 14K Math and English responses from rural and urban schools. Using this dataset, we evaluate VLMs and LLMs for grading and personalized feedback, finding that GPT-4o and DeepSeek (671B) perform competitively with teacher scores, with DeepSeek producing more contextually relevant feedback. This work lays the foundation for AI-powered assessment in low-resource and underrepresented educational contexts.
@@ Acknowledgments
Reserved due to double-blinds.
@@
Appendix A Example of Assessment Items and Rubrics
Example assessment items for English and Math are shown below. The questions were designed in accordance with the Indonesian curriculum to ensure that students can understand the questions given.
  • Math
    (Multiple choice) Pecahan mana yang paling besar? (Which fraction is the largest?) A. 1/9 B. 1/8 C. 1/7 D. 1/6
    (Essay) Terdapat empat buah pecahan: 3/10, 8/10, 1/10, 6/10. Urutkan pecahan-pecahan tersebut dari yang paling kecil ke yang paling besar (There are four fractions: 3/10, 8/10, 1/10, 6/10. Sort the fractions from smallest to largest)
  • English
    (Multiple choice) Look at the picture and choose the correct word for it [Picture is provided] A. bathroom; B. living room; C. Dining room
    (Essay) Look at the picture. [Picture is provided] Write a text, consisting of information: How many rooms are there? (write 1 sentence) What are the rooms? (write 1 sentence) How is the position of the rooms? (Write at least 5 sentences)
Moreover, we also provide an example of a rubric used to score the essay here. The rubric was developed by a teacher with over 10 years of classroom experience, ensuring it is accurate and aligned with the assessment criteria.
  • Math (written in Indonesian)
    Score 0: Jawaban ada dan benar, tetapi tanpa alasan (The answer is correct, but without a reason)
    Score 1: Jawaban ada dan benar, hanya menyebutkan 1/10 lebih besar dari 1/100. (The answer is correct, it just states that 1/10 is greater than 1/100.)
    Score 2: Jawaban ada dan benar, hanya menyebutkan 10 lebih kecil dari 100. (The answer is correct, it only states that 10 is less than 100.)
    Score 3: Jawaban ada dan benar, menyebutkan karena pecahan tersebut adalah pecahan dengan pembilang satu sehingga semakin kecil penyebut, semakin besar pecahannya. (The answer is correct, stating that the fraction is a fraction with a numerator of one, so the smaller the denominator, the larger the fraction.)
  • English (written in English)
    Score 0: The response does not meet the above criteria (e.g., incorrect sentence structure, incorrect use of prepositions, or no response).
    Score 1: The response is an incomplete sentence, but uses the correct preposition of place. Example: "bedroom under the living room."
    Score 2: The student’s response is a complete sentence but has 1 or 2 punctuation errors. Example: "The bedroom is under living room."
    Score 3: The student’s response is a complete sentence with correct grammar and punctuation. Example: "The bedroom is under the living room."
Appendix B Hyperarameter Setup
@@
Appendix C Prompts List
@@
Appendix D Human Evaluation Guideline on Personalised Feedback
@@
Appendix E Qwen-VL-72B OCR Results and The Decision of Using GPT-4o
For OCR, we decided to use GPT-4o instead of another open-source LLM, as we had already conducted a comparison between GPT-4o and Qwen-VL-72B across two subjects: English and Math. The results for Qwen-VL-72B are shown in Table 5, with an average English score of 67.5% and a Math score of 35.1%, which are both slightly lower than GPT-4o’s performance (English: 68%, Math: 36.5%). Therefore, we chose GPT-4o for the OCR tasks.
English Math
Area EM(M) EM(S) RL(E) EM(M) EM(S) RL(E)
Urban 80.9 67.2 59.3 59.9 23.3 20.5
Rural 73.8 60.2 60.2 60.1 21.8 23.8
All 78.5 64.7 59.3 61.0 22.8 21.6
Table 5: OCR-based performance (Qwen-VL-72B) across Urban, Rural, and All settings for English and Math: EM = exact match, RL = ROUGE-L F1, M = multiple choice, S = Short Essay, E = Essay.
Appendix F Additional Data Details: Students’ Average Scores
For completeness, we also present the dataset information, including the average scores for each subject and question type. The students’ average scores, manually graded by humans, are shown in Table 6. Based on these scores, we calculate the mean absolute error (MAE) for each model, as presented in Table 1, with English and Math evaluated separately according to each model’s grading performance.
English Math
Area M S E M S E
Urban 85.27 63.29 30.90 52.00 67.43 61.36
Rural 72.29 41.21 26.60 34.80 63.38 66.71
All 80.53 55.22 30.90 52.00 67.43 61.36
Table 6: The students’ average scores, manually graded by humans. M = multiple choice, S = Short Essay, E = Essay.
Appendix G Example Qualitative Analysis
We also present examples of students’ assessments, specifically from rural areas, as shown in Figures 5 and 6. Based on the OCR results, certain patterns indicate that the OCR model (e.g., GPT-4o) struggles to interpret some student handwriting, even though it performs well on multiple-choice questions.
In the case of the math subject, the OCR model occasionally misinterprets numerals. For example, reading a 9 in the denominator as an 8, or incorrectly extracting 1/3 as 1/5. For the English subject, most errors occur in the essay section. The OCR model sometimes misreads words, such as interpreting “bathroom” as “bedroom.” Additionally, the model often fails to predict the answer to the final English question correctly because some students write more responses than the paper allows (a common occurrence in real-life classroom settings), which confuses the OCR system.
Figure 5: Example of a student’s math assessment answer sheet from a rural area. The student’s identity has been obscured for privacy.
Figure 6: Example of a student’s English assessment answer sheet from a rural area. The student’s identity has been obscured for privacy.